OpenStudy (sleepyjess):

Convert 210° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed. \(\dfrac{7\pi}{12}\) \(\dfrac{7\pi}{3}\) \(\dfrac{7\pi}{5}\) \(\dfrac{7\pi}{6}\)

OpenStudy (sleepyjess):

I got \(\dfrac{210\pi}{80}\) simplified to \(\dfrac{21\pi}{8}\)

OpenStudy (mayankdevnani):

Remember :- When you convert degrees into radians ,we use this suppose `x` be a degree :- \[\large \bf Radian=\frac{\pi}{180}\times x\]

OpenStudy (mayankdevnani):

so,for 210 degrees :- \[\large \bf Radians=\frac{\pi}{180} \times 210\]

OpenStudy (mayankdevnani):

now simplify it !

OpenStudy (mayankdevnani):

@sleepyjess i hope you understand !

OpenStudy (sleepyjess):

OOOHHHHHHH i put 80 and not 180

OpenStudy (mayankdevnani):

oh!! that's your mistake though

OpenStudy (mayankdevnani):

now tell me what you get after simplfying it ??

OpenStudy (sleepyjess):

\(\dfrac{7\pi}{6}\)

OpenStudy (mayankdevnani):

great !! good job !

OpenStudy (sleepyjess):

Thank you for helping me realize my stupid mistake.

OpenStudy (mayankdevnani):

your welcome :)

Nnesha (nnesha):

this way :) is use this formula  now cross multiply